What is the (b) in the equation 2=(3/2)(3)+b using the formula y=mx+b
Answer by jim_thompson5910(35256)

2=(3/2)(3)+b

2=9/2+b

2-9/2=b

4/2-9/2=b

-5/2=b

So the answer is b = -5/2